OpenStudy (anonymous):

Being driven crazy by radical expressions. Hoping someone will understand what I am trying to say here. 3root-64/27

OpenStudy (anonymous):

\[\sqrt[3]{-64/27}\] what about this question?

OpenStudy (anonymous):

Thats it! It says to evaluate if possible

OpenStudy (nowhereman):

\[\sqrt[3]{-\frac {64}{ 27}} = -\frac{ \sqrt[3]{64}} {\sqrt[3]{27}} = -\frac 4 3\]

OpenStudy (amistre64):

the radical can be split up so that one radical covers each number

OpenStudy (amistre64):

nowhere beat me to it :) good job

OpenStudy (anonymous):

-4/3 can also be -1.33333333...

OpenStudy (anonymous):

Thanks nowhereman!! These things are killing me!

OpenStudy (nowhereman):

Well, you look alive to me :-D